OpenStudy (anonymous):

im getting better and better at these questions but i get messed up when there is a negative. solve the following problem -14+17= A.60 B.-60 C.26 D.-26 remember im asking for help not answers

jimthompson5910 (jim_thompson5910):

-14+17 is the same as 17 + (-14) which is also the same as 17 - 14

OpenStudy (anonymous):

which is 26? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

No. Oddly enough, none of your answer choices are the correct answer. Is there a typo?

OpenStudy (anonymous):

oh sorry there is a typo didnt even relize that its -43+17 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

-43+17 can be thought of as 43-17 but the final result is negative (since -43 is larger in magnitude and it's negative)

OpenStudy (anonymous):

-43+14=26

OpenStudy (anonymous):

-26

jimthompson5910 (jim_thompson5910):

I think you meant to say -43 + 17 = -26

OpenStudy (anonymous):

yes i did wow with all these typos!

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @jim_thompson5910 -43+17 can be thought of as 43-17 but the final result is negative (since -43 is larger in magnitude and it's negative) \(\color{blue}{\text{End of Quote}}\) 17 -43

OpenStudy (anonymous):

here is another one i think i already figured out but can you check it? @jim_thompson5910 -20-54+4= and i got -18 @Nnesha

OpenStudy (anonymous):

sorry i got -81 omg

OpenStudy (anonymous):

wow i have to type this again it -20-65+4 and i got -81 @jim_thompson5910 @Nnesha sorry with all the typos

jimthompson5910 (jim_thompson5910):

-20-65+4 = -81 is correct